# Knowledge-to-SQL: Enhancing SQL Generation with Data Expert LLM

### v2

@@ 3.3 Preference Learning with PLDBF
Typically, preference fine-tuning is employed subsequent to supervised fine-tuning for further enhancement. In our scenario, we adopt a direct preference optimization (DPO) Rafailov et al. (2023) algorithm in our preference learning framework. For each preference data $(I_{l},K_{w},K_{l})\in\mathcal{D}$, the objective of this process can be formulated as:
@@
$=-E_{\pi}[\log\sigma(\beta R(K_{w})-\beta R(K_{l}))],$
where $E_{\pi}=\mathbb{E}_{(I_{l},K_{w},K_{l})\sim\mathcal{D}}$ and $R(K)=\log(\pi^{DPO}(K|I_{l})/\pi^{SFT}(K|I_{l}))$, R is the reward implicitly defined by the target language model $\pi^{DPO}$ and reference model $\pi^{SFT}$. We get the DPO enhanced model $\pi^{DPO}$, which can generate a knowledge assist accurate database execution and SQL contributing.
@@ 3.4 Knowledge-to-SQL with DELLM
Finally, the user query, the database schema, and the relevant table are given during the testing time. With preference-enhanced DELLM $\pi^{DPO}$, we generate knowledge and combine it with the query, schema, and instruction to assist an off-the-shelf text-to-SQL model for SQL generation. We enumerate each query and the corresponding schema and get the predicted SQL as the results.
@@ 4 Experiments
In this section, we will empirically evaluate our proposed framework, i.e., Knowledge-to-SQL. After introducing the experimental setups, the evaluation discussions are divided into five parts:
  • Main Result: The aim of our framework is to assist a text-to-SQL model in generating accurate SQL. We will compare the originally predicted SQL with the SQL generated using our acquired knowledge, assisting various models or methods on different benchmarks.
  • Ablation Study: We will experiment with the ablation study to verify the efficiency of our framework. By predicting SQL using the knowledge generated by variations of our proposed framework, we will discuss the influence of different components.
  • Comprehensive Assistance Evaluation: We firstly look deep into what query type (difficulty) the knowledge assists the SQL generation. Then, we compare the improvement of our generated knowledge to the improvement of the ground-truth (human expert annotated) knowledge. And conclude the comparison based on the two results above.
@@
  • Statistic Analysis: To look deep into the impact of DELLM from the data statistics aspect, we calculate the ratio of potential influences brought by the DELLM’s generating knowledge and discuss the impact.
@@ 4.1 Experimental Setup
Our experiments are conducted on two widely used dataset BIRD Li et al. (2023b) and Spider Yu et al. (2018). BIRD benchmark was released recently, which annotated high-quality text-to-SQL instances with 95 databases on a large scale. BIRD is a leading benchmark focusing on massive and real database content, first introducing external knowledge reasoning between natural language questions and database content. As the evaluation metric, BIRD introduces a new metric verifying the balance of efficiency and accuracy during the execution. Spider benchmark is assessed frequently to evaluate the performance of text-to-SQL across multiple databases, which include 200 distinct databases and 138 domains. The test set of these two datasets is not publicly available, so we evaluate the efficacy of our method on the accessible development set.
@@
To fairly compare the results with previous approaches, we follow the metric as previous work Gao et al. (2023); Wang et al. (2023). We consider two metrics, execution accuracy and valid efficiency score, to evaluate our generated knowledge on prompting text-to-SQL models. We consider the metrics 1) Execution Accuracy (EX), which is the ratio of questions that the execution results of the predicted SQLs match exactly with those of the ground-truth SQLs, compared to the total number of questions. 2) Valid Efficiency Score (VES), designed to evaluate the efficiency of SQLs generated by text-to-SQL models. It emphasizes that the ”valid SQLs” refer to predicted SQLs with execution results identical to those of the ground-truth SQLs.
Implementations.
We utilize LLaMA-2-13b Touvron et al. (2023) with official configuration as our backbone model for knowledge generation. The learning rate is set as 5e-05, selected from the interval [1e-05, 1e-04]. The other hyper-parameters are discussed in Appendix A. Noting that the Spider dataset does not have annotated knowledge as training samples, we utilize BIRD knowledge to train the model and evaluate the main performance in both the BIRD and Spider datasets. The evaluation of the Spider dataset aims to verify our effectiveness in cross-domain database scenarios. The ablation study and further discussion are evaluated on the BIRD dataset.
Baselines.
We introduce the comparative baselines on the BIRD dataset first. Specifically, to verify the assistance on open-source model, we utilize 1) T5-3B Raffel et al. (2020), using the fine-tuning-based method and taking the knowledge as fine-tuning input. Then, for the off-the-shelf models, we adopt: 2) GPT-3.5-Turbo and 3) GPT-4 OpenAI (2023), a widely-use powerful proprietary model with zero-shot text-to-SQL prompt for SQL generation. 4) Claude-2 Anthropic (2023), another popular proprietary model. For the recently introduced prompt engineering techniques, we compare: 5) GPT-3.5-Turbo + COT Li et al. (2023b), a simple chain-of-thoughts prompt engineering on text-to-SQL. 6) DAIL-SQL Gao et al. (2023), encoding structure knowledge and selects few-shot examples based on similarity matching. 7) MAC-SQL Wang et al. (2023), a novel LLM-based multi-agent collaborative framework designed for the text-to-SQL task. On Spider dataset, as introduced above, the evaluation is conducted on a cross-domain database. We adopt proprietary models GPT-3.5-Turbo and GPT-4 as verification.
4.2 Main Results
@@
Table 2: Experimental results for text-to-SQL on different benchmarks with and without knowledge generated by our method. The number in the bracket denotes the improvement of DELLM’s knowledge compared to the without knowledge baseline.
Table 2 gives the results as the verification of our method. The number in the brackets represents the improvement of the metric by prompting the LLMs with knowledge generated by our method.
@@
The knowledge generated by DELLM obtain promising results on both metrics in assisting different models/prompting techniques in the text-to-SQL task, which obtain around 5% improvement on EX and 6% in VES around all comparative baselines. Specifically, 1) The knowledge significantly improves the model’s performance with simple prompting techniques. We owe this performance to the preference enhancement of the database execution and ground-truth SQL contribution feedback, which enable the model DELLM to generate helpful knowledge to introduce database schema and the user query, assisting SQL generation. The straightforward prompt brings a great challenge to the schema and query understanding ability of LLMs, especially when the query is challenging. DELLM brings a substantial improvement to the simple prompt. 2) With prompting techniques, the knowledge also gains promising results. The prompting technique obtains an improvement compared to the simple prompt by focusing on providing related examples or task decomposing, which is a various angle from knowledge assisting, to enhance SQL generation. This means that when facing the scenario without annotated knowledge, DELLM can potentially improve the state-of-the-art performance from a knowledge-assisting angle. Although their prompting technique obtains great improvement, there is still space for promotion with helpful knowledge.
@@
The database scenario of Spider is less challenging. The performance on the same metrics is overall better, as we can see from the results. Similar to BIRD: 1) The knowledge substantially improves the performance of proprietary models. 2) The improvement brought by the knowledge is significantly lower than the improvement on the BIRD dataset. We owe this result to the different complexity and characteristics of the database structure in the two datasets and the difficulty of the user query. The query in the Spider dataset may not be as challenging as in the BIRD. As the number of queries that can be helped decreases, the improvement declines respectively.
4.3 Abalation Study
Table 3 presents the results of an ablation study for our framework in the BIRD dev set with proprietary LLM GPT-4. The table lists different variations of our framework, including excluding database execution feedback (db feedback) or ground-truth SQL contributing feedback (sql feedback) and also removing the table reading process. The column presents the evaluation metrics.
@@ 4.3 Abalation Study
Table 3: Ablation study on variations of our proposed framework.
As we can see in the table, removing any components leads to decreased performance on both metrics. When removing the database feedback, the performance decline the most. Assisting the predicted SQL in database execution is the priority in generating helpful knowledge. Overall, the ablation study indicates that each component of our method plays an irreplaceable role in achieving good performance, as their removal resulted in decreased performance as intuitive expectation.
4.4 Comprehensive Assistance Evaluation
To get a comprehensive conclusion, we discuss SQL generation on different difficulty levels and compare the knowledge generated by DELLM to the ground-truth knowledge. Table 4 shows the results on the BIRD dev set, decomposing different difficulty levels of the user query. In the table, D and E represent the knowledge generated by our proposed DELLM (D) and the ground-truth knowledge annotated by human experts (E).
@@ 4.4 Comprehensive Assistance Evaluation
Table 4: The execution accuracy (EX) for queries with different difficulty. Simp. denotes Simple, Mod. denotes Moderata, and Chall denotes Challenging.
From the results: 1) There is a gap between the DELLM-generated knowledge and the expert-annotated one on all difficulty levels. The expert-annotation knowledge surpasses 8.67% and 8.73% to DELLM on execution accuracy, leaving a great gap for improvement. 2) The generated knowledge mainly works on simple and moderate queries. The reason is that LLMs have difficulty understanding the challenging query and schema, making it hard to utilize corresponding generated knowledge.
@@ 4.5 Performance on Partial Training Data
We select partial data in the BIRD annotated knowledge training set to train our model. And discuss the performance of that scenario. We conduct the experiments on GPT-4, and the results are shown as the improvement with the generated knowledge by DELLM on different ratios of training data.
@@
Figure 3: Improvement to GPT-4 on different metrics with knowledge generated by DELLM with different ratios of training data.
As shown in Figure 3, where the x-axis represents the ratio of training data, the y-axis is the value of the metric’s improvement. In practice, with a limited budget, we can hire the database expert to annotate about 60% of training data to gain around 77% performance for DELLM knowledge generating in text-to-SQL improvement.
4.6 Statistic Analysis
We identify four influences that the DELLM’s knowledge may bring: 1) Assistance: denote the original predicted SQL is not valid, but becomes predicted valid with the knowledge. 2) Misleading: denote the original predicted SQL is valid, but invalid influenced by the knowledge. 3) Inoperative: represent the original invalid SQL still be invalid with the knowledge. 4) Sustainable: represent the predicted SQL is valid and still remains valid with the knowledge. The experiment is conducted using GPT-4 on execution accuracy.
Figure 4: Different influence of DELLM brings on GPT-4’s performance.
From Figure 4, we can see the DELLM brings around 5% assistance, and there is potentially 0.5% misleading of its generated knowledge. The inoperative knowledge leaves about 60% in DELLM and also around 53% in the expert annotation. One potential optimization to text-to-SQL is improving the performance of the knowledge understanding.
@@ 5 Conclusion
In conclusion, our study highlights the significance of knowledge generation in enhancing the performance of large language models (LLMs) for text-to-SQL tasks. By proposing a novel framework that leverages database information, execution feedback, and ground-truth comparisons, we address the challenges of bridging the semantic gap between user queries and database schemas. Through empirical validation on Bird and Spider datasets, our approach demonstrates substantial improvements in execution accuracy for models like GPT-4, underscoring its efficacy in advancing text-to-SQL research and fostering innovations in natural language processing and data mining.
@@ 6 Limitations
The knowledge generated by our approach relies on the in-context learning and text-to-SQL capability of the LLMs when we cooperate with these LLMs performing text-to-SQL tasks. With the help of DELLM, online LLMs such as GPT4 and GPT3.5, which only have publicly available APIs, have obtained significant text-to-SQL performance. However, the in-context learning capability and text-to-SQL capability of open-source medium-scale local LLMs still have a large gap comprised of GPT4. With the help of our generated knowledge, even their text-to-SQL performance has achieved a higher than 50 percent improvement in performance, which is still not as good as accessing the GPT4 API. This may limit the application of DELLM in offline deployment situations.
7 Ethics Statement

### v3

@@ 3.3 Preference Learning with PLDBF
Typically, preference fine-tuning is employed subsequent to SFT for further refinement. In our scenario, the preference learning framework utilizes a direct preference optimization (DPO) Rafailov et al. (2023) algorithm. For each preference pair $(I_{l},K_{w},K_{l})\in\mathcal{D}$, the objective function of this process can be formulated as:
@@
$=-\mathbb{E}_{\pi}[\log\sigma(\beta R(K_{w})-\beta R(K_{l}))],$
specifically, $\mathbb{E}_{\pi}=\mathbb{E}_{(I_{l},K_{w},K_{l})\sim\mathcal{D}}$ and $R(K)=\log(\pi^{DPO}(K|I_{l})/\pi^{SFT}(K|I_{l}))$. where $R(\cdot)$ is the reward implicitly defined by the target model $\pi^{DPO}$ and reference model $\pi^{SFT}$. Then, we obtain the DPO refined model $\pi^{DPO}$, which can generate expert knowledge to assist in accurate database execution and contribute to precise SQL generation.
@@ 3.4 Knowledge-to-SQL with DELLM
Finally, the user question, the database schema, and the relevant table are given during the testing phase. With PL-refined DELLM $\pi^{DPO}$, we collect the generated expert knowledge and combine it with the question, schema, and task instruction to assist an off-the-shelf text-to-SQL model for SQL generation. We enumerate each question and the corresponding schema and get the predicted SQL query as the result.
@@ 4 Experiments
In this section, we will empirically evaluate our proposed Knowledge-to-SQL framework. After introducing the experimental setups, the experimental results are discussed in five parts:
  • Main Result: The purpose of our framework is to assist a text-to-SQL model in generating accurate SQL. We compare the originally predicted SQL with the predicted SQL incorporating the generated expert knowledge; the comparative performance is evaluated on various models or methods on different benchmarks.
  • Ablation Study: We conduct the ablation studies to verify the efficiency and robustness of our proposed framework. By predicting SQL using the knowledge generated by variations of DELLM, we discuss the influence of different modules.
  • Comprehensive Evaluation: We first look deep into how the generated knowledge assists different types (difficulties) of the question. Then, we compare the improvement brought by the generated knowledge and the ground-truth knowledge. Then conclude the comparative results based on the two results above.
@@
  • Statistical Analysis: To visualize the influence of DELLM from a data statistical perspective, we calculate the ratio of various influences brought by incorporating the generated knowledge and provide corresponding discussions and analyses.
@@ 4.1 Experimental Setup
Our experiments are conducted on two widely recognized dataset BIRD Li et al. (2023b) and Spider Yu et al. (2018). BIRD benchmark was released most recently, which annotated high-quality text-to-SQL instances with 95 databases on a large scale. BIRD is a leading benchmark focusing on massive and real database content, first introducing knowledge reasoning between natural language questions and database content. As the evaluation metric, BIRD introduces a new metric verifying the balance of efficiency and accuracy during the execution. The knowledge introduced by BIRD is annotated by human experts who are native speakers of English with degrees above the bachelor’s level and have database-related knowledge. Spider benchmark is assessed frequently to evaluate the performance of text-to-SQL across multiple databases, which include 200 distinct databases and 138 domains. Since the test set of these two datasets is not publicly available, we evaluate our method’s efficacy on the accessible development (dev) set.
@@
To make a fair comparison, we follow the metric as previous work Gao et al. (2024); Wang et al. (2023) for evaluation. We consider two metrics: 1) Execution Accuracy (EX), which is the ratio of questions that the execution results of the predicted SQL queries match exactly with those of the ground-truth SQL queries, compared to the total number of questions. 2) Valid Efficiency Score (VES), designed to evaluate the efficiency of SQL generated by text-to-SQL systems. The VES metric considers both the efficiency and accuracy of execution results and incorporates running time for a more comprehensive evaluation.
@@
Table 2: Experimental results for text-to-SQL on different benchmarks with and without knowledge generated by our proposed DELLM. The number in the bracket denotes the improvement in execution accuracy (EX) and valid efficiency score (VES) brought by DELLM’s knowledge compared to the baseline performance without knowledge.
Implementations.
We select the widely-used LLaMA-2-13b Touvron et al. (2023) with official configuration as our backbone model for knowledge generation. The learning rate is set as 5e-05, selected from the interval [1e-05, 1e-04]. The semantic similarity calculation of the table reading process is implemented by Faiss [^2] https://github.com/facebookresearch/faiss/ Douze et al. (2024). The other hyper-parameters are discussed in the Appendix A. Noting that the Spider dataset does not have annotated knowledge samples, we utilize the human-annotated knowledge in BIRD as ground-truth knowledge to train the model and evaluate the main performance in both the BIRD and Spider datasets. The evaluation of the Spider dataset aims to verify our effectiveness in cross-domain databases. The ablation study and further discussions are conducted on the BIRD dataset.
Baselines.
We select the comparative baselines on the BIRD dataset based on the official benchmark page [^3] https://bird-bench.github.io/. Specifically, to verify the effectiveness of DELLM on the open-source model, we utilize 1) T5-3B Raffel et al. (2020), using the fine-tuning-based method that incorporates the knowledge as fine-tuning input. Then, for the off-the-shelf models, we utilize 2) GPT-3.5-Turbo and 3) GPT-4 OpenAI (2023), a widely-used powerful proprietary model with zero-shot text-to-SQL prompt for SQL generation; 4) Claude-2 Anthropic (2023), another well-recognized proprietary model. For the up-to-date prompt engineering techniques, we compare 5) GPT-3.5-Turbo + CoT Li et al. (2023b), a simple CoT prompt engineering on text-to-SQL; 6) DAIL-SQL Gao et al. (2024), encoding structure knowledge and selects few-shot instances based on similarity matching; 7) MAC-SQL Wang et al. (2023), a novel LLM-based multi-agent collaborative framework designed for the text-to-SQL task. On the Spider dataset, as introduced above, the evaluation is conducted as a cross-domain verification. We utilize proprietary models GPT-3.5-Turbo and GPT-4 for validation.
4.2 Main Results
Table 2 gives the experimental results. The number in the brackets represents the improvement of the metric by prompting the LLMs with knowledge generated by DELLM.
@@ 4.2 Main Results
The knowledge generated by DELLM obtained promising results on both metrics in assisting different models/prompting techniques in the text-to-SQL task, which obtained around 5% improvement on EX and 6% in VES around all comparative baselines. Specifically, 1) The knowledge significantly improves the model’s performance with simple prompting. We owe this phenomenon to the PL-refinement based on the database execution and ground-truth SQL contribution feedback, which enable DELLM to generate more helpful knowledge to assist SQL generation. The straightforward prompt challenges the LLMs’ capability for question and schema understanding, especially when the question is challenging. DELLM generates the corresponding expert knowledge, reducing the difficulty of understanding, which leads to substantial improvement. 2) When assisting prompting techniques, the knowledge also achieves promising results. The prompting technique surpasses the simple prompt by focusing on providing high-quality few-shot instances or decomposition, which is a various angle from knowledge that assists in enhancing SQL generation. This means that DELLM can potentially improve the state-of-the-art performance from a knowledge-assisting angle when facing the scenario without annotated knowledge. Although the well-designed techniques achieve solid performance, there is still space for improvement in generating helpful knowledge.
@@
The database scenario of Spider is less challenging. The performance on the same metrics is overall better, as we can see from the results. Similar to BIRD, 1) The knowledge substantially improves the performance of proprietary models. 2) The improvement brought by the knowledge is significantly lower than the BIRD dataset. We owe this result to the divergent complexity and characteristics of the database structure and the different difficulties of the user question. The question in the Spider dataset may not be as challenging as in the BIRD. As the number of questions that need to be assisted decreases, the improvement declines respectively.
4.3 Ablation Study
Table 3 presents the results of the ablation study for our framework in the BIRD dev set with proprietary LLM GPT-4. The table lists different variations of DELLM, including excluding database execution feedback (db feedback) or ground-truth SQL contributing feedback (sql feedback) and also removing the table reading process. The column presents the evaluation metrics.
@@ 4.3 Ablation Study
Table 3: Ablation study on variations of DELLM.
As we can see in Table 3, removing any components leads to a performance decrease on both metrics. The performance declined the most when excluding the database feedback, indicating that assisting the predicted SQL in database execution is the priority in generating helpful knowledge. Overall, the ablation study demonstrates that each component of our method plays an irreplaceable role in achieving good performance, as their removal decreased the performance as intuitively expected.
We also discuss the choice of PL algorithm. Apart from the utilized direct preference optimization (DPO) Rafailov et al. (2023) in our framework, we evaluate a variation of DELLM, which is PL-refined by proximal policy optimization (PPO) Schulman et al. (2017).
Models EX VES
GPT-4 33.25 $\pm$ 0.61 35.92 $\pm$ 1.03
+ $\text{DELLM}_{PPO}$ 35.28 $\pm$ 1.18 40.03 $\pm$ 1.81
+ $\text{DELLM}_{DPO}$ 37.94 $\pm$ 0.57 42.15 $\pm$ 0.95
Table 4: Ablation study on the utilized PL algorithms.
Given the results in Table 4, we can conclude: 1) PPO works for DELLM, which can also generate helpful knowledge to improve the performance of the text-to-SQL model. 2) DPO outperforms PPO, one potential reason is that PPO relies on a reward model to give numeric rewards for each generated knowledge. Since the knowledge generation task is non-deterministic, even if the provided knowledge is correct, the downstream SQL generation may be wrong. In this case, the DPO algorithm with implicit rewards performs better.
4.4 Comprehensive Evaluation
To reach a comprehensive conclusion, we discuss SQL generation at different difficulty levels and compare the knowledge generated by DELLM to the ground-truth knowledge. Table 5 shows the results on the BIRD dev set, decomposing different difficulty levels of the user question. In the table, D and E represent the knowledge generated by our proposed DELLM (D) and the ground-truth knowledge annotated by human experts (E).
@@ 4.4 Comprehensive Evaluation
Table 5: The execution accuracy (EX) for questions with different difficulty. Simp. denotes Simple, Mod. denotes Moderata, and Chall denotes Challenging.
From the results, 1) There is a gap between the DELLM-generated knowledge and the human expert-annotated one on all difficulty levels. The human-annotated knowledge surpasses 8.67% and 8.73% to DELLM on execution accuracy, leaving a great gap for improvement. 2) The generated knowledge mainly works on simple and moderate questions. One reason is that LLMs have difficulty understanding the challenging question and schema, making it hard to utilize corresponding generated knowledge. Another reason is that challenging questions only occupy a small portion of the annotations (around 9%), leading to a data imbalance during the training process.
@@ 4.5 Performance on Partial Training Data
We select partial annotated knowledge in the BIRD training set to train the DELLM, then discuss the performance of that scenario. We conducted the experiments using GPT-4, and the results are shown as the improvement with the generated knowledge by DELLM on different ratios of training data.
@@
Figure 3: Improvement to GPT-4 on different metrics with DELLM on different ratios of training data.
As shown in Figure 3, where the x-axis represents the ratio of training data, the y-axis is the value of the metric’s improvement. In practice, with a limited budget, we can hire human experts to annotate about 60% of training data to obtain around 77% performance for expert knowledge generation of DELLM.
4.6 Statistical Analysis
We identify four types of influences that the DELLM’s knowledge may bring: 1) Assistance, where the original predicted SQL is not correct but becomes correct with the help of the knowledge. 2) Misleading, where the original predicted SQL is correct, but becomes incorrect due to the influence of the knowledge. 3) Inoperative, where the original incorrect SQL remains incorrect despite the knowledge. 4) Sustainable, where the originally predicted SQL is correct and remains correct with the knowledge. The statistics are obtained based on the experiment result using GPT-4 on the BIRD dev set.
Figure 4: Different influences of DELLM bring on GPT-4’s performance on the BIRD dev set.
From Figure 4, the DELLM brings around 5% assistance, and there is potentially 0.5% misleading of its generated knowledge. The inoperative knowledge leaves about 60% in DELLM and also around 53% in the expert annotation. One potential optimization to text-to-SQL is improving the effectiveness of incorporating expert knowledge.
@@ 5 Conclusion
In conclusion, our study highlights the significance of knowledge generation in enhancing the performance of LLMs for text-to-SQL tasks. By proposing a novel framework that leverages database content, execution feedback, and comparisons with ground-truth SQL, we address the challenges of bridging the knowledge gap between user questions and database schema. Through extensive experiments on BIRD and Spider datasets, our approach brings substantial improvements in execution accuracy and valid efficiency scores for models like GPT-4, underscoring its efficacy in advancing text-to-SQL research and fostering innovations in natural language processing and data mining.
@@ 6 Limitations
Two primary limitations are acknowledged in this study. Firstly, the performance of DELLM is mostly evaluated using proprietary models. However, the in-context learning and natural language understanding capability of open-source medium-scale local LLMs still have a large gap compared to proprietary LLMs. Although with the assistance of generated knowledge, their performance significantly improved, the overall performance is not good enough for practical use. This may limit the application of DELLM in offline deployment situations. Secondly, the generalizability of the proposed framework to real-world scenarios remains to be determined, as it has only been evaluated on standard benchmarks. The complexity, diversity, and potential noise in real-world database environments calls for further verification to confirm the framework’s effectiveness in practical applications.
Ethics Statement
@@
Acknowledgements
This work was supported in part by the Innovation and Technology Commission of HKSAR and the Ministry of Science and Technology of China under the Innovation and Technology Fund - Mainland-Hong Kong Joint Funding Scheme (MHP/012/21), and in part by the Innovation and Technology Commission of HKSAR under the Research Talent Hub for ITF projects (Pih/054/23).